### v1

# A Quantile Nelson-Siegel model

Abstract
A widespread approach to modelling the interaction between macroeconomic variables and the yield curve relies on three latent factors usually interpreted as the level, slope, and curvature (Diebold et al., 2006). This approach is inherently focused on the conditional mean of the yields and postulates a dynamic linear model where the latent factors smoothly change over time.
However, periods of deep crisis, such as the Great Recession and the recent pandemic, have highlighted the importance of statistical models that account for asymmetric shocks and are able to forecast the tails of a variable’s distribution. A new version of the dynamic three-factor model is proposed to address this issue based on quantile regressions. The novel approach leverages the potential of quantile regression to model the entire (conditional) distribution of the yields instead of restricting to its mean.
An application to US data from the 1970s shows the significant heterogeneity of the interactions between financial and macroeconomic variables across different quantiles. Moreover, an out-of-sample forecasting exercise showcases the proposed method’s advantages in predicting the yield distribution tails compared to the standard conditional mean model. Finally, by inspecting the posterior distribution of the three factors during the recent major crises, new evidence is found that supports the greater and longer-lasting negative impact of the great recession on the yields compared to the COVID-19 pandemic.
Keywords: Bayesian inference; Nelson-Siegel model; Quantile regression; Time-varying parameters; Yield curve.
1 Introduction
The ongoing global challenges stemming from the COVID-19 pandemic and recent geopolitical conflicts have prompted a renewed focus on inflation and monetary policy. Of particular significance is the information embedded in the term structure of interest rates, offering valuable insights into the evolution of these economic factors. The slope of the yield curve has gained considerable recognition as a key predictor of economic activity, exerting a substantial influence on investors’ decision-making processes.
This significance is underscored by research from Benzoni et al. (2018) and Haubrich (2021), which highlights a strong correlation between an inversion of the yield curve and the onset of recessions. For instance, in May 2019, the yield curve inverted almost a year before the commencement of the most recent recession in March 2020. More recently, Fonseca et al. (2023) emphasised the unprecedented inversion in the risk-free yield curves in the United States due to the rapid rise in short-term interest rates since 2022. They pointed out that “recession probability models based solely on the yield curve slope currently point to elevated odds of recession in the euro area and the United States in one year’s time” thereby stressing the critical importance of dealing with and studying the yield curve in contemporary macroeconomic analysis.
Transitioning to a macroeconomic perspective, significant time variations have been shown to influence the dynamics of US inflation and real activity (Mumtaz and Surico, 2009). Mönch (2012) studied the evolution of macroeconomic variables by using level, slope, and curvature factors, finding evidence of the informativeness of the curvature factor about the future evolution of the yield curve. Moving to factor analysis, Coroneo et al. (2016) provided evidence that some common factors drive the dynamics of macroeconomic variables and government bonds. More recently, Fernandes and Vieira (2019) predicted the yield curve using different forward-looking macroeconomic variables and highlighted the importance of these variables in forecasting the yield curve.
In the context of modelling the term structure of interest rates, the Nelson-Siegel (NS) model, introduced by Nelson and Siegel (1987), is the most popular empirical framework employed within the literature. Diebold and Li (2006) generalised this model by introducing smooth dynamics for the latent factors, resulting in the dynamic NS (DNS) model, which outperforms the standard NS approach in an out-of-sample forecasting context. However, these approaches do not account for the potential impact of macroeconomic variables on the yield curve. To address this limitation, Diebold et al. (2006) and Coroneo et al. (2016) analysed the dynamic linkages between yield curve factors and macroeconomic factors in the US. Koopman and van der Wel (2013) found evidence of interdependence between macroeconomic and latent factors in the US term structure, thus supporting the use of macroeconomic variables in combination with financial yield curve factors. Continuing along this line of research, Bianchi et al. (2009) applied a macro-factors augmented DNS model with time-varying parameters to UK data. Recently, Han et al. (2021) introduced a time-varying decay parameter for the arbitrage-free NS model and demonstrated that the relative factor loading reaches its peak just before starting the decline right before a recession.
It is worth emphasising that the literature on the term structure and the analysis of the relationship between yield curve factors and macroeconomic variables so far has been concerned with conditional mean models. However, the in-sample and forecasting performances of this approach are only sufficient in relation to the mean of the yields, as the method is not suited to investigate the complete distribution of the yields. To address the limitations of standard conditional mean Nelson-Siegel models and capture the complete picture of the conditional distribution of the yields, we propose a novel quantile regression Nelson-Siegel model with time-varying parameters (TVP-QR-NS) model.
We contribute to this literature by proposing a quantile version of the Nelson-Siegel model for the yield curve. Specifically, we extend the Diebold et al. (2006) ’Yields-Macro’ model to a quantile framework. Our approach assumes that the conditional $\tau$-quantile of the yields is given by the Nelson-Siegel three-factor specification, where the factors are (smoothly) time-varying. Quantile regression (QR, see Koenker and Bassett, 1978) does not assume a parametric likelihood for the conditional distribution of the response variable. Unlike standard models based on the conditional mean and relying on symmetric second-moment dynamics, our novel framework offers robust modelling of the conditional quantiles, thus enabling a comprehensive investigation of the entire conditional distribution. Moreover, by considering multiple quantile levels, we can follow the method proposed by Mitchell et al. (2022) and combine the quantiles to obtain the distributional yield curve at each point in time, as shown in Figure 1. This permits the investigation of dynamic asymmetry within the distribution of the term structure of interest rates. This feature is particularly crucial when undertaking the modelling and forecasting of interest rate term structures, as emphasised in our empirical application.
Figure 1: Distributional term structure of the interest rate for June and December 2022.
Inference is performed within a Bayesian framework by leveraging the connection between a likelihood formed by the asymmetric Laplace distribution (Kotz et al., 2001) and quantile regression provided by Yu and Moyeed (2001), and Petrella and Raponi (2019).
The main feature of the proposed TVP-QR-NS model is the coupling of quantile regression with time-varying three-factor Nelson-Siegel specification. In addition, to cope with the potentially high dimensionality of the latent dynamic factors, we use a precision sampler (Chan and Jeliazkov, 2009; Chan et al., 2020) that allows us to sample the entire path of the factors jointly without any loop. Overall, this combination provides a simple framework and computationally efficient method to investigate and forecast the term structure at every quantile level of interest.
The proposed method is then applied to investigate and forecast the US term structure of interest rates. Specifically, an in-sample analysis highlights the advantages of the quantile approach compared to the standard conditional mean Nelson-Siegel model. In particular, we are able to obtain the entire distribution of the three factors at different points in time, thus enabling a more detailed analysis of the effects of the recent crises (the Great Recession and the COVID-19 pandemic) on the term structure. Interestingly, these results, which could not be obtained using a conditional mean approach, reveal substantial differences between the impacts of the financial crisis and COVID-19. The former shifted leftwards the distribution of the short-term interest rate without meaningful impact on the long-term. Conversely, the pandemic had an initial significant adverse effect on both the short- and long-term rates, but the distribution in 2022 shifted rightwards compared to the pre-COVID one.
The advantages of the proposed TVP-QR-NS model are better highlighted in terms of tail risk yield forecasting. Compared to the conditional mean model, our approach provides better out-of-sample forecasts in both tails of the yield distribution, as measured by the quantile score. The difference is particularly striking on the left tail and consistent over time.
The remainder of this article is organised as follows. Section 2 introduces and describes the novel framework, then Section 3 presents the Bayesian approach to inference. Section 4 details the empirical application. Finally, Section 5 concludes.
2 Methodology
2.1 Nelson-Siegel Yield Curve model
The standard Nelson-Siegel (NS) model of the yield curve assumes that
$y_{t}(m)=\beta_{1}+\beta_{2}\left(\frac{1-\exp^{-\lambda m}}{\lambda m}\right)+\beta_{3}\left(\frac{1-\exp^{-\lambda m}}{\lambda m}-\exp^{-\lambda m}\right),$ (1)
where $m$ denotes the time to maturity; that is, given a set of parameters, the yield curve is a deterministic function of maturity. Diebold and Li (2006) pioneered the stochastic modelling of the yield curve and suggested the linear model
$y_{t}(m)=\beta_{1t}+\beta_{2t}\left(\frac{1-\exp^{-\lambda m}}{\lambda m}\right)+\beta_{3t}\left(\frac{1-\exp^{-\lambda m}}{\lambda m}-\exp^{-\lambda m}\right)+\epsilon_{t}(m)$ (2)
where $\epsilon_{t}(m)\sim\mathcal{N}(0,\sigma^{2}(m))$ is an independent normal innovation. Therefore, given a set of time-varying parameters, the expected value of the curve is a deterministic function of maturity.
2.2 A Time-Varying Parameters Quantile NS Model
Assume that $y_{t}(m)\in\mathcal{Y}\subset\mathds{R}$ for all maturities $m\in\mathds{R}_{+}$, and let $F_{t}(\cdot;m)$ denote the cumulative distribution function of $y_{t}(m)$. We propose a new approach that alleviates the parametric assumptions of $y_{t}(m)$’s by modelling the quantile function of $F$,
$Q_{t}^{\tau}(m)=\inf\left\{x\in\mathcal{Y}:F_{t}(x;m)\geq\tau\right\}$
for each quantile level $\tau\in(0,1)$ as follows
$Q_{t}^{\tau}(m)=\beta_{1t}(\tau)+\beta_{2t}(\tau)\Big{(}\frac{1-e^{-\lambda^{\tau}m}}{\lambda^{\tau}m}\Big{)}+\beta_{3t}(\tau)\Big{(}\frac{1-e^{-\lambda^{\tau}m}}{\lambda^{\tau}m}-e^{-\lambda^{\tau}m}\Big{)},$ (3)
where $m$ is the maturity, $\beta_{t1}^{\tau}$, $\beta_{t2}^{\tau}$, and $\beta_{t3}^{\tau}$ denote the (unobserved) level, slope and curvature factors, and $\lambda^{\tau}\in\mathds{R}_{+}$ is a decay parameter. In the following, we shall assume $\lambda^{\tau}=\lambda$. Thus, by directly modelling the quantile function, we allow the yield curve factors to behave differently across the quantile levels.
Consider a series of observations at specific maturities $m\in\{1,\ldots,M\}$ and denote with $\mathbf{y}_{t}=(y_{t}(m_{1}),\ldots,y_{t}(m_{M}))^{\prime}$ the vector of yields at different maturity. Let $\mathbf{x}(m;\lambda)=\big{(}1,\frac{1-e^{-\lambda m}}{\lambda m},\frac{1-e^{-\lambda m}}{\lambda m}-e^{-\lambda m}\big{)}^{\prime}$, and $\mathbf{X}(\lambda)=\left[\mathbf{x}(m_{1};\lambda),\mathbf{x}(m_{2};\lambda),\ldots,\mathbf{x}(m_{M};\lambda)\right]^{\prime}\in\mathds{R}^{M\times 3}$. Then, the conditional quantile model can be written as
$Q^{\tau}_{t}=\mathbf{X}(\lambda)\bm{\beta}_{t}(\tau)$ (4)
such that $Q^{\tau}_{t}=(Q^{\tau}_{t}(m_{1}),\ldots,Q^{\tau}_{t}(m_{M}))^{\prime}$ and $\bm{\beta}_{t}(\tau)=(\beta_{1t}(\tau),\beta_{2t}(\tau),\beta_{3t}(\tau))^{\prime}$. Moreover, at each quantile level $\tau$, the unobserved factors are modelled jointly with a vector of macroeconomic (observed) variables $\mathbf{z}_{t}\in\mathds{R}^{r}$. Specifically, the vector of time-varying (unobserved) yield curve and (observed) macroeconomic factors $\bm{\phi}_{t}=(\bm{\beta}_{t}^{\prime}(\tau),\mathbf{z}_{t}^{\prime})^{\prime}$ is assumed to follow a stationary VAR(1) process
$\bm{\phi}_{t}=\bm{\gamma}_{0}+\Gamma\bm{\phi}_{t-1}+\bm{\eta}_{t},\qquad\bm{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\Omega).$ (5)
Differently from Coroneo et al. (2016) that proposed an approach related to ours in the context of a conditional mean model for the yields, we are modelling the conditional quantile of $y_{t}(m)$ and rely on an alternative specification of the interactions between the yield and the macroeconomic factors.
To design a regression model for the yield such that the conditional $\tau$-quantile corresponds to eq. (3), we adopt a likelihood-based approach and rely on the properties of the asymmetric Laplace distribution (Kotz et al., 2001). In particular, for a given quantile $\tau$, we consider a regression model where the innovation follows an asymmetric Laplace distribution $\epsilon_{t}(m)\sim\mathcal{AL}(\mu,\sigma,\tau)$, with location $\mu\in\mathds{R}$, scale $\sigma>0$, skewness parameter $\tau\in(0,1)$, with density
$p(x|\mu,\sigma,\tau)=\frac{\tau(1-\tau)}{\sigma}\exp\Big{\{}-\rho_{\tau}\Big{(}\frac{x-\mu}{\sigma}\Big{)}\Big{\}},$
where $\rho_{\tau}(x)=x(\tau-\mathbb{I}(x\leq\tau))$ is the check-loss function. Therefore, combining the conditional quantile in eq. (3) with the asymmetric Laplace innovation results in
$y_{t}(m)=\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau).$ (6)
It is important to highlight that the asymmetric Laplace distribution admits a stochastic representation as a location-scale mixture of Gaussian (Kotz et al., 2001), thus allowing to rewrite eq. (6) as
$y_{t}(m)\stackrel{{\scriptstyle d}}{{=}}\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}(\tau)+w_{m,t}\sigma\theta_{\tau,1}+\sqrt{w_{m,t}}\sigma\theta_{\tau,2}u_{t}(m),\qquad u_{t}(m)\sim\mathcal{N}(0,1),$ (7)
where $w_{m,t}\sim\mathcal{E}xp(1)$ is exponential with rate one. To establish an equivalence between the mode of the asymmetric Laplace distribution and the optimisation of the standard quantile regression problem in Koenker and Bassett (1978), the parameters $\theta_{\tau,1},\theta_{\tau,2}$ are constrained to the following values (Kotz et al., 2001):
$\theta_{\tau,1}=\frac{1-2\tau}{\tau(1-\tau)},\qquad\theta_{\tau,2}=\sqrt{\frac{2}{\tau(1-\tau)}}.$ (8)
Throughout the article, we follow the existing literature and fix the decay parameter at $\lambda=0.0609$ (e.g., see Diebold and Li, 2006). Moreover, the scale of the innovation is set to $\sigma=1.0$, as in the original quantile regression model (Yu and Moyeed, 2001).
The combination of eq. (5) and (6) provides the state-space form of the proposed time-varying parameters quantile Nelson-Siegel model (TVP-QR-NS) as
$\begin{split}y_{t}(m)&=\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau)\\ \bm{\phi}_{t}&=\bm{\gamma}_{0}+\Gamma\bm{\phi}_{t-1}+\bm{\eta}_{t},\qquad\bm{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\Omega),\end{split}$ (9)
where we recall the fixed values $\lambda=0.0609$ and $\sigma=1.0$. Denoting with $\mathbf{Y}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{T})^{\prime}\in\mathds{R}^{T\times M}$ the matrix of observations, the joint density of the observables and auxiliary variables (conditional on the time-varying factors) is
$p(\mathbf{Y},\mathbf{w}|\bm{\beta})$ $=\prod_{m=1}^{M}\mathcal{E}xp(w_{m,t}|1)\times\mathcal{N}\big{(}y_{t}(m)|\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+w_{m,t}\sigma\theta_{\tau,1},\>w_{m,t}\sigma^{2}\theta_{\tau,2}^{2}\big{)}.$ (10)
3 Estimation
This section is devoted to the prior description and the derivation of the precision sampler algorithm.
3.1 Prior Specification
We consider a multivariate Gaussian prior for the initial value of the time-varying factors:
$\bm{\beta}_{0}(\tau)\sim\mathcal{N}(\underline{\bm{\mu}}_{\beta},\underline{\Sigma}_{\beta}).$ (11)
By leveraging the mixture representation of the asymmetric Laplace distribution, we obtain a conditionally linear Gaussian state space for the latent dynamic vector $\bm{\beta}_{t}$. Thus, we sample from the posterior full conditional distribution of $\bm{\beta}_{t}$ directly in a single loop, using a precision sampler (e.g., see Chan and Jeliazkov, 2009; Chan et al., 2020) as described in the following subsection.
Let us define $e_{m,t}=y_{t}(m)-\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}$. Then, the posterior full conditional distribution of the auxiliary variable $w_{t}$ is obtained as
$w_{m,t}|y_{t}(m),\bm{\beta}_{t}$ $\propto w_{m,t}^{-\frac{1}{2}}\exp\Big{\{}-\frac{1}{2}\Big{[}\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}}w_{m,t}+2w_{m,t}+w_{m,t}^{-1}\frac{e_{m,t}^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}\Big{]}\Big{\}},$ (12)
which is the kernel of the generalized inverse Gaussian distribution $\text{GiG}(\overline{p}_{w},\overline{a}_{w},\overline{b}_{w,m,t})$ with
$\overline{p}_{w}=1-\frac{1}{2},\qquad\overline{a}_{w}$ $=2+\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}},\qquad\overline{b}_{w,m,t}=\frac{e_{m,t}^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}.$
Finally, we specify a Gaussian prior for the vectorised coefficient matrix of the state equation, $\bm{\gamma}=(\bm{\gamma}_{0}^{\prime},\operatorname{vec}(\Gamma)^{\prime})^{\prime}$, and an inverse Wishart prior for the innovation covariance, that is
$\bm{\gamma}\sim\mathcal{N}(\underline{\bm{\mu}}_{\Gamma},\underline{\Sigma}_{\Gamma}),\qquad\Omega\sim\mathcal{IW}(\underline{v},\underline{V}).$ (13)
As the prior distributions are conjugate, the full conditional posteriors are straightforward to obtain as
$\bm{\gamma}|\bm{\phi},\Omega\sim\mathcal{N}(\overline{\bm{\mu}}_{\Gamma},\overline{\Sigma}_{\Gamma}),\qquad\Omega|\bm{\phi},\gamma_{0},\Gamma\sim\mathcal{IW}(\overline{v},\overline{V}).$ (14)
The derivation of these full conditional posteriors is standard (e.g, see Chan, 2020).
3.2 The precision sampler
While the conditional distribution of the parameters given the data and the set of unobserved latent factors is relatively standard, the most significant computational costs arise from sampling the large number of latent factors, denoted as $\bm{\beta}_{t}(\tau)$.
Let us start by rewriting the measurement equation in (9) in compact form by stacking all observations together. Specifically, we define the stacked vectors $\mathbf{y}=(\mathbf{y}_{1}^{\prime},\ldots,\mathbf{y}_{T}^{\prime})^{\prime}$, $\bm{\beta}=(\bm{\beta}_{1}(\tau)^{\prime},\ldots,\bm{\beta}_{T}(\tau)^{\prime})^{\prime}$, $\mathbf{w}=(w_{1},\ldots,w_{T})^{\prime}$, and $\mathbf{u}=(u_{1},\ldots,u_{T})^{\prime}$, we have
$\mathbf{y}=\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\bm{\beta}+\theta_{\tau,1}\sigma\mathbf{w}+\mathbf{u},$ (15)
where $\mathbf{u}\sim\mathcal{N}(\mathbf{0}_{TM},\sigma^{2}\theta_{\tau,2}\Omega_{1})$ with $\Omega_{1}=\operatorname{diag}(\mathbf{w})$. Similarly, we define $\bm{\phi}=(\bm{\phi}_{1}^{\prime},\ldots,\bm{\phi}_{T}^{\prime})^{\prime}$ and rewrite the state equation in a matrix form as
$\mathbf{H}\bm{\phi}=\tilde{\bm{\gamma}}+\bm{\eta},\qquad\bm{\eta}\sim\mathcal{N}\left(\mathbf{0}_{T(r+3)},\Omega\otimes\mathbf{I}_{T}\right),$ (16)
where
$\mathbf{H}=\begin{pmatrix}\mathbf{I}_{r+3}&\mathbf{0}_{r+3}&\mathbf{0}_{r+3}&\mathbf{0}_{r+3}&...&\mathbf{0}_{r+3}&\mathbf{0}_{r+3}\\ \mathbf{0}_{r+3}&-\Gamma&\mathbf{I}_{r+3}&\mathbf{0}_{r+3}&...&\mathbf{0}_{r+3}&\mathbf{0}_{r+3}\\ ...&...&...&...&...&...&...\\ ...&...&...&...&...&\mathbf{I}_{r+3}&\mathbf{0}_{r+3}\\ ...&...&...&...&...&-\Gamma&\mathbf{I}_{r+3}\end{pmatrix},\qquad\tilde{\bm{\gamma}}=\begin{pmatrix}\bm{\gamma}_{0}\\ \bm{\gamma}_{0}\\ ...\\ \bm{\gamma}_{0}\\ \bm{\gamma}_{0}\end{pmatrix}.$
Note that the unobserved factors are placed first in the VAR vector such that
$\bm{\phi}=S_{1}\bm{\beta}+S_{2}\mathbf{u},$
where
$S_{1}=\mathbf{I}_{T}\otimes(\mathbf{I}_{3},\mathbf{0}_{3\times r})^{\prime},\quad S_{2}=\mathbf{I}_{T}\otimes(\mathbf{0}_{3\times r},\mathbf{I}_{r})^{\prime}.$
Thus, we obtain the conditional posterior of the time-varying yield curve factors as $\bm{\beta}|\mathbf{Y},\mathbf{w},\Gamma,\bm{\gamma}_{0},\Omega\propto\mathcal{N}\left(\bm{\mu}_{b},K_{b}^{-1}\right)$, where
$K_{b}$ $=\frac{1}{\sigma^{2}\theta_{\tau,2}}\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)^{\prime}\Omega_{1}^{-1}\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)+(\mathbf{H}S_{1})^{\prime}\mathbf{I}_{T}\otimes\Omega^{-1}(\mathbf{H}S_{1}),$
$\bm{\mu}_{b}$ $=K_{b}^{-1}\left(\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)^{\prime}\Omega_{1}^{-1}(\mathbf{y}-\theta_{\tau_{1}}\sigma\mathbf{w})+S_{1}\mathbf{H}^{\prime}\mathbf{I}_{T}\otimes\Omega^{-1}\left(\bm{\gamma}-\mathbf{H}S_{2}z\right)\right).$
4 Empirical Application
This section is devoted to studying the effectiveness of the proposed TVP-QR-NS model for zero-coupon US treasury yields. In particular, we split the following section into an in-sample analysis of the yield curve factor and a forecasting exercise.
4.1 Data
Our empirical investigation employs zero-coupon US Treasury yields as constructed by Liu and Wu (2021) through a non-parametric kernel smoothing methodology. The application of this novel yield curve dataset by Liu and Wu (2021) is demonstrated to offer a more precise depiction of the underlying data when compared to the alternative metric presented by Gürkaynak et al. (2007). Following the methodology established by Diebold et al. (2006), our analysis encompasses zero-coupon US Treasury maturities spanning 3, 6, 9, 12, 15, 18, 21, 24, 30, 36, 48, 60, 72, 84, 96, 108, and 120 months. Additionally, our investigation includes key monthly macroeconomic indicators — Industrial Production, CPI Inflation, and the Fed Funds Rate — in line with the framework suggested by Diebold et al. (2006). These macroeconomic variables are chosen to capture the dynamic aspects of the macroeconomy and are all sourced from the US FRED database. Both Industrial Production and CPI Inflation were transformed to log-differenced growth rates. Our study period extends from November 1971 to December 2022.
4.2 In-Sample Results
Previous studies, including Diebold and Li (2006); Diebold et al. (2006, 2008); Mumtaz and Surico (2009); Mönch (2012), focused on the yield curve and its effects within a conditional mean framework. Our study is distinct from the former as it is the first to explore the yield curve from a quantile or distributional perspective. This section presents our empirical in-sample results. The first subsection details the posterior estimates of yield curve factors for each quantile level. The second subsection analyses the distributional features of short- and long-term interest rates during recent crises (the Great Recession and the COVID-19 pandemic). Lastly, the third subsection investigates the interaction between yield curve factors and the macroeconomy from a quantile perspective.
4.2.1 Quantile Yield Curve Factor Estimates
Figures 2, 3, and 4 depict the posterior mean estimates of time-varying yield curve factors at five quantiles (10%, 25%, 50%, 75%, 90%), providing insights into their dynamic evolution. A consistent finding across all graphs is the presence of substantial temporal variation. Beginning with the level factor, interpreted as the long-term interest rate (i.e., yield as maturity approaches infinity), a steady decline in this rate is discernible across all quantiles since the oil price crisis of the 1980s. Figure 2 reveals a distinct pattern: following both the dot-com recession in the early 2000s and the Great Recession of 2008-09, the posterior estimates for the 90th and median quantiles closely align, while those for the 10th quantile diverge. Conversely, after the COVID-19 recession, the opposite trend emerged, with the 10th and median quantiles exhibiting proximity while the 90th quantile diverged. This evidence suggests that post the early 2000s recession and the Great Recession, the distribution of the long-term interest rate skewed leftwards. In contrast, after the COVID-19 recession, the distribution skewed rightwards, likely attributed to the pronounced inflationary pressures experienced in the aftermath of the COVID-19 period.
Figure 2: Posterior mean estimates of the level factor for each of the five quantiles (10%, 25%, 50%, 75%, 90%). NBER recession periods in grey shades.
Turning to the slope and curvature of the yield curve in Figures 3 and 4, the degree of uncertainty associated with these factors varies over time. Notably, the curvature factor exhibits higher volatility across all quantiles than its slope counterpart. [^1] Notice that crossing of the estimated conditional quantiles for the curvature factor is not harmful since it does not have a direct link to one particular yield. A distinct pattern is evident before 1985, where the posterior estimates for the slope factor across all five quantiles are closely clustered, indicating a period of low uncertainty. In contrast, during the oil price crisis of the 1980s, the curvature factor displayed substantial dispersion across quantiles.
Figure 3: Posterior mean estimates of the slope factor for each of the five quantiles (10%, 25%, 50%, 75%, 90%). NBER recession periods in grey shades.
Post-recessionary periods following the early 1990s, the 2000s, and the Great Recession of 2008-09 reveal elevated uncertainty in both slope and curvature factors. Specifically, convergence is observed between the median and the 90th quantile, with the 10th quantile exhibiting a discernible deviation from this convergent trajectory. In the aftermath of the COVID-19 recession, a reverse trajectory emerges, with the 90th quantile diverging from the 10th and median quantiles. The associated uncertainty surrounding both factors appears comparatively subdued compared to the preceding recessions of the 2000s. Both the early 2000s recession and the Great Recession were attributed to financial crises. Our empirical findings suggest a potential connection between periods following financial crises and increased uncertainty in the slope and curvature of the yield curve.
Figure 4: Posterior mean estimates of the curvature factor for each of the five quantiles (10%, 25%, 50%, 75%, 90%). NBER recession periods in grey shades.
It is worthwhile mentioning that these results could not be obtained using a conditional mean approach, which is instead able to produce estimates of the mean yield factors over time. Conversely, the proposed quantile approach enables deeper and more detailed analyses of the term structure while maintaining a low computational cost.
4.2.2 The Distributional Features of the Short- and Long-Term Interest Rate
The proposed novel framework provides a distinct advantage by facilitating the investigation of distributional characteristics in both short- and long-term interest rates over time. As mentioned earlier, the long-term interest rate is identified as the level factor, while the short-term interest rate is a composite of the level and slope factors. We adhere to the methodology outlined by Mitchell et al. (2022) to construct the distributions of these interest rates. Employing an equal weighting scheme, we aggregate all MCMC draws across five quantiles and apply the Epanechnikov kernel, as proposed by Gaglianone and Lima (2012).
The previous analysis revealed an increase in the degree of dispersion among quantiles during post-recessionary periods, particularly in the 2000s. Expanding on this finding, we investigate whether the distributional features of short- and long-term interest rates exhibit temporal variations across the two most recent recessions. Figure 5 illustrates the distributions of both short- and long-term interest rates for selected months preceding and following the Great Recession and the COVID-19 recession.
Figure 5: distribution of the short-term and long-term (level) factors during the Great Recession (left) and the COVID-19 pandemic (right).
Focusing on the short-term interest rate first, discernible evidence of a pronounced distributional shift is observed across the three selected months for both recessions. Notably, during the Great Recession (September 2009), a manifestation of multi-modality in the distribution is apparent. Additionally, after the COVID-19 recession (December 2020), the distribution became less symmetric and more negatively skewed to the left. Turning to the long-term interest rate, the evidence of distribution shifts is more prominent during the COVID-19 recession than during the Great Recession. The long-term interest rate distribution exhibits increased leftward skewness during the Great Recession (September 2009) and the subsequent period (December 2009). Conversely, during the COVID-19 recession, the distribution becomes more positively skewed to the right. Therefore, our results highlight that both short- and long-term interest rate distributions display heightened skewness and asymmetric shapes during recessionary periods. Thus highlighting the importance of using quantile models instead of standard Nelson-Siegel models that are not able to detect the left and right tails of the distribution.
4.2.3 The Interaction between the Yield Curve and Macroeconomy - Quantile VAR Parameter estimates
The proposed framework, as delineated in eq. (9), introduces an avenue for interaction between yield curve factors and macroeconomic variables by incorporating a VAR-like state equation. The estimated off-diagonal VAR coefficients for various quantile levels are presented in Figure 6, shedding light on the directional effects between yield curve factors and macroeconomic variables.
Panel A
Panel B
Figure 6: Posterior mean and credible intervals for the VAR coefficients across different quantile levels.
Panel A of the figure conveys the directional impact of yield curve factors on macroeconomic variables, while Panel B reports the converse scenario. The patterns of influence of each yield factor on the three macroeconomic variables exhibit remarkable similarity. Precisely, evidence of a U-shaped effect across quantiles is discerned for industrial production and inflation, with an inverse relationship observed for the Fed funds rate. Conversely, the impact of macroeconomic variables on the yield curve factor demonstrates higher heterogeneity. Industrial production exerts a significant positive effect on both level and curvature, whereas its impact on the slope is mixed. The effects of inflation exhibit qualitative similarity, albeit with a smaller magnitude on the level and a higher impact on the slope. In contrast, the Fed funds rate positively influences all yield curve factors. Analysing the relationship with quantile levels, the effects on the slope exhibit a U-shaped pattern for industrial production and inflation, while the Fed funds rate demonstrates an inverse U-shaped relationship. Meanwhile, the relationship with the level remains constant across quantiles. In summary, our findings provide substantial evidence that the interaction between yield curve factors and the macroeconomy significantly varies across quantiles.
4.3 Forecasting Yields Tail Risk
We conduct a pseudo-out-of-sample forecasting exercise to assess the utility of our novel quantile Nelson-Siegel approach, comparing it against the corresponding standard Nelson-Siegel conditional mean counterpart for forecasting tail risks in zero-coupon yields. The initial holdout period spans from November 1971 to January 2000, with a subsequent forecast evaluation period extending from February 2000 to December 2022 (comprising 264 evaluation periods). We generate 1st, 6th, and 12th step-ahead tail risk forecasts and employ the quantile score (QS) (see Giacomini and Komunjer, 2005; Carriero et al., 2022) as the forecast evaluation metric to gauge the performance of the two models. Following Gneiting and Ranjan (2011), we define the QS as:
$QS_{\tau,i,t}=(y_{t}(m)-\mathcal{Q}_{\tau,i,t})-(\tau-\mathbb{I}\{y_{t}(m)\leq\mathcal{Q}_{\tau,i,t}\}).$ (17)
Here, $QS_{\tau,i,t}$ denotes the predictive quantile of the $m$-th yield, and $\mathbb{I}\{y_{t}(m)\leq\mathcal{Q}_{\tau,i,t}\}$ takes the value 1 if the realised value is at or below the predictive quantile and 0 otherwise. We evaluate the QS in the upper and lower tails by setting $\tau=0.9$ and $\tau=0.1$, respectively.
Table 1 presents the 10th and 90th percentiles of QS for the out-of-sample forecasting exercise. The top panel of Table 1 displays the average QS of our proposed approach relative to their corresponding conditional mean counterparts for each of the 17 yields across the evaluation period. A quantile score less than 1 implies that our proposed approach outperforms the corresponding conditional mean counterpart. Notably, our proposed approach consistently produces superior tail forecasts for all 17 yields and across all forecast horizons compared to their corresponding conditional mean counterparts. We further assess the significance of our forecasting results through a Diebold-Mariano test, with results presented in the bottom panel of Table 1, clearly indicating the statistical significance of our forecasting results in favour of our proposed approach.
Additionally, we depict the average raw QS across the 17 yields for both models over time in Figure 7. The top and bottom panels show the 10th and 90th percentiles of QS, respectively. Three key features emerge from this figure. First, reinforcing the results in Table 1, our proposed quantile Nelson-Siegel model consistently outperforms their corresponding conditional mean counterparts across all horizons and quantile levels, with the evidence being stronger for 1-step ahead forecasts. Second, the 1-step ahead QS (for both 10th and 90th percentiles) for the conditional mean model exhibits significantly more instability over time than our quantile approach. Finally, we find strong evidence in favour of the proposed time-varying parameters quantile Nelson-Siegel model for forecasting the left tail of the yield distribution, suggesting its ability to more appropriately capture the lower bound of the yield curve compared to standard conventional models.
Figure 7: Quantile Score over time for the 1-, 6-, and 12-steps ahead forecast horizons, at quantile levels 10% (top plots) and 90% (bottom plots), for the QR-NS model (blue) and the standard NS model (red). Shaded areas represent NBER recession periods.
Table 1: Forecasting results for the quantile NS relative to standard NS model, together with Diebold-Mariano test. Out-of-sample evaluation period: 2000 - 2022.
Ratio of quantile scores (quantile NS over standard NS)
10% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 0.28 0.34 0.39 0.43 0.44 0.45 0.45 0.45 0.43 0.42 0.41 0.42 0.45 0.50 0.54 0.60 0.65 0.45
6-step ahead 0.61 0.63 0.64 0.65 0.65 0.65 0.64 0.64 0.63 0.61 0.59 0.58 0.57 0.57 0.57 0.58 0.59 0.61
12-step ahead 0.67 0.68 0.69 0.70 0.70 0.71 0.71 0.70 0.70 0.69 0.67 0.66 0.65 0.63 0.63 0.62 0.62 0.67
90% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 0.47 0.55 0.56 0.55 0.53 0.51 0.50 0.49 0.49 0.49 0.50 0.53 0.54 0.56 0.57 0.57 0.56 0.53
6-step ahead 0.98 0.98 0.96 0.93 0.90 0.88 0.86 0.84 0.80 0.77 0.73 0.71 0.69 0.69 0.69 0.69 0.68 0.81
12-step ahead 0.96 0.96 0.95 0.93 0.92 0.91 0.89 0.88 0.86 0.84 0.82 0.80 0.79 0.78 0.77 0.76 0.75 0.86
Diebold-Mariano test, significance level: 1 (10%), 2 (5%), 3 (1%)
10% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
6-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
12-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
90% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
6-step ahead 0.00 0.00 2.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 2.59
12-step ahead 1.00 1.00 2.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 2.71
5 Conclusions
In this study, we introduce a novel time-varying parameters quantile Nelson-Siegel (TVP-QR-NS) model within a Bayesian framework and apply it to an empirical analysis of the US term structure of interest rates. Our in-sample findings reveal an increased dispersion in yield curve factors across quantiles in the post-recession period. Additionally, temporal variations in the distribution of short- and long-term interest rates are observed during the two most recent recessions. Furthermore, evidence indicates significant fluctuations in the interaction between yield curve factors and the macroeconomy across quantiles. Finally, through an out-of-sample forecasting exercise, we demonstrate the superior performance of our proposed approach in tail risk forecasting across all yields compared to its standard conditional mean counterpart.